# Optimal Trudinger-Moser inequalities on complete noncompact Riemannian manifolds: Revisit of the argument from the local inequalities to global ones

### v2

@@ Abstract.
The main purpose of this short note, on the one hand, is to clarify some part of the proof of Theorem 1.3 in [10] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma. Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls (or the charts). This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set and thus rigorize the proof of Theorem 1.3 in [10]. This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
@@ 1. Introduction
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded, i.e. $\textit{Ric}\geq\lambda g$ for some constant $\lambda\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. Then for every $\tau>0$, there exists a constant $C=C(n,\tau,M)$ such that
@@
In the proof of Theorem 1.1 some clarifications need to be made about the uniformity of the upper bound of the Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the Trudinger-Moser inequality on geodesic balls within the manifolds (still going through the same proof of Theorem 3.1 in [10] on the geodesic balls or simply applying Theorem 3.1 in [10] on geodesic balls where the Green functions have uniform estimates) together with a Gromov’s covering lemma (similar to the proof of subcritical case in [10]). The proof of Theorem 3.1 in [10] remains valid for geodesic balls with uniform constants for the Trudinger-Moser inequalities for such balls independent of the location of the balls.
The main result of this note is that the Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius and Ricci lower bound have uniform estimates independent of the locations of the balls. This will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set. More precisely, in the proof of Theorem 1.3 in [10], the level set $\Omega(u)$ can be covered by a finite number of small geodesic balls (or harmonic coordinate charts) and the number of such balls (or charts) is uniformly bounded. Therefore, the same level argument from local inequalities to the global ones in [10] are justified.
@@ 2. Some preliminaries in differential geometry
On the other hand, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates. This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
@@
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in\Omega$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
@@
Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$ and we have $W^{2,p}$ uniform bounds.
3. Green’s function estimate
@@
where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 near diagonal and vanishes when $d(x,y)\geq\textit{inj}(M,g)$. $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
@@
where $C=C(n,\lambda,Q,\lambda,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
@@
From the asymptotic behavior near singularity, we need to prove the following lemma:
@@
(3.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z)$
It is well known that in polar coordinates
$dV=r^{n-1}\left(1-\frac{\textit{Ric}(\theta)}{3}r^{2}+o(r^{2})\right)d\theta dr.$
When $\delta$ is small, providing Ricci curvature lower bounded, the first term of the right hand side,
@@
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $C(\lambda,\delta)\omega_{n-1}\rho^{\alpha+\beta-n},$
Similarly we can estimate the second term of the right hand side of (3.3). For the third term, we have
@@
(3.7) $\leq$ $C_{n}C(\lambda,\delta)\omega_{n-1}\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$

### v3

@@ Abstract.
The main purpose of this short note, on the one hand, is to rigorize some part of the proof of Theorem 1.3 in [11] using the level set argument in a simple way, and on the other hand, to give an alternative argument from local inequalities to global ones by using Gromov’s covering lemma. This result holds on Riemannian manifolds with positive injective radius and Ricci lower bound as stated in [11]. Moreover, we show that the Green functions for the Dirichlet problem on any small geodesic balls (or harmonic coordinate charts) on Riemannian manifolds with additional condition that the sectional curvature is bounded from above have uniform estimates independent of the locations of the balls (or the charts). This will give another proof on such manifolds the critical Trudinger-Moser inequality. This latter result of Green’s function estimates is of its independent interest and can be found useful elsewhere.
@@ 1. Introduction
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded, i.e. $\textit{Ric}\geq\lambda g$ for some constant $\lambda\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. Then there exists a positive constant $\epsilon=\epsilon(\lambda,i,n)=O(\frac{1}{r_{H}})$ in the reverse order of the harmonic radius $r_{H}$ such that for every $\tau>\epsilon$, there exists a constant $C=C(n,\tau,M)$ such that
@@
In the proof of Theorem 1.1 needs to be rigorized about the uniformity of the upper bound of the critical Trudinger-Moser inequality in terms of the bounded domains of the level sets involved. In this note, we will then simply give another simple proof by only establishing the uniformity of the bounds of the critical Trudinger-Moser inequality on geodesic balls within the manifolds by modifying the ideas developed in the earlier work by Lam, Lu and Zhang [10]. In fact, we will achieve this by using a scaling argument and the asymptotic estimates for the supremum of the subcritical Trudinger-Moser functional established in [10]. As a consequence of this, we can prove the uniform boundedness of the critical Trudinger-Moser inequality on the level set of the functions under consideration in [11]. This thus rigorizes the proof of Theorem 1.1 (namely, Theorem 1.3 in [11]) by covering the level set by the geodesic balls. Therefore, the same level argument from local inequalities to the global ones used in [11] are justified. We make use of the equivalence between the subcritical and critical Moser-Trudinger inequalties on geodesic balls. This idea was originally in [10] and already used in [11] to derive subcritical inequalities from critical inequalities. Meanwhile in this note, we will prove in the opposite direction, i.e. from subcritical inequalities to critical inequalities.
The second main result of this note is to establish that Green functions for the Dirichlet problem on any small geodesic balls on Riemannian manifolds with positive injective radius, Ricci curvature lower bounded and sectional curvature upper bounded have uniform estimates independent of the locations of the balls. We believe that this result is of its independent interest. More precisely,
Theorem 1.2.
Let $(M,g)$ be a complete noncompact Riemannian manifold whose Ricci curvature is lower bounded and sectional curvature is upper bounded, i.e. $\textit{Ric}\geq\lambda g$ and $K\leq a^{2}$ for some constant $\lambda,a\in\mathbb{R}$. Moreover, we assume that its injectivity radius has a lower bound, i.e. $\textit{inj}(M,g)\geq i>0$ for some constant $i>0$. For any $Q>1$ and $0<\alpha<1$, denote the geodesic ball as $B_{\delta}(x)$ with $\delta\leq r_{H}(n,i,\lambda,\alpha)$, where $r_{H}(n,i,\lambda,\alpha)$ is the harmonic radius. Then the Dirichlet Green’s function over $B_{\delta}(x)$ satisfies the following estimate:
$|\nabla_{y}G(x,y)|\leq\frac{1}{\omega_{n-1}}d(x,y)^{1-n}(1+Cd(x,y)),$
where $C=C(n,\lambda,Q,\delta)$.
As a byproduct, this will allow us to conclude that the Trudinger-Moser inequality on any level set is uniformly bounded independent of the location of the level set, within an extra sectional curvature assumption by using the uniform estimates of the Green functions.
@@ 2. Some preliminaries in differential geometry
[…]
$\begin{cases}\Delta x^{i}(Q)=0,\ \ \ \textit{for}\ Q\in B_{\rho}(P);\\ x^{i}(P)=0,\frac{\partial x^{i}}{\partial\widetilde{x}^{j}}(P)=\delta^{i}_{j};\\ x^{i}(Q)=\widetilde{x}^{i}(Q)\ \ \ \textit{for}\ Q\in\partial B_{\rho}(P).\end{cases}$
Its solution $x^{i}$ is called harmonic coordinate for $B_{\rho}(P)$. Moreover, lower bounds for both Ricci curvature and injectivity radius guarantee the existence of $C^{0,\alpha}$ harmonic coordinates. This is due to Anderson-Cheeger [1] (see also Hebey [6]) and to be precise:
@@
Let $\alpha\in(0,1)$, $Q>1$, $\delta>0$ and $(M,g)$ satisfy $\textit{Ric}\geq\lambda g$ and $\textit{inj}(M,g)\geq i>0$. Then there exists a positive constant $C=C(n,Q,\alpha,\delta,i,\lambda)$ such that for any $x\in M$, the $C^{0,\alpha}$ harmonic radius $r_{H}(Q,0,\alpha)(x)\geq C$.
@@
Similar embedding also appears for the coordinate functions $x^{i},i=1,\dots,n$. We have $W^{2,p}$ , and hence $C^{1,\alpha}$ uniform bounds.
3. A path from subcritical inequalities to critical ones: Proof of Theorem 1.1
This section gives a proof of Theorem 1.1 by establishing uniform critical Moser-Trudinger inequalities on geodesic balls from subcritical ones. By a covering lemma, we can establish the global ones either by using the level set argument as done in [11] or the covering lemma.
We first recall some notations.
Notation.
Let $(M,g)$ be an $n$-dimensional Riemannian manifold, $n\geq 2$. Fix $\tau>0$ and define
$\left\lVert u\right\rVert_{1,\tau}:=\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\tau\left\lVert u\right\rVert_{L^{n}(M)}.$
Let
$\phi_{n}(t):=e^{t}-\sum_{k=0}^{n-2}\frac{t^{k}}{k!},\qquad\alpha_{n}:=n\Bigl(\frac{n\pi^{n/2}}{\Gamma(\frac{n}{2}+1)}\Bigr)^{\!\!\frac{1}{n-1}}=n\,\omega_{n-1}^{1/(n-1)}.$
We first quote a result from [10]. The following theorem provides the lower and upper bounds asymptotically for the supremum of the subcritical Trudinger-Moser inequality in $\mathbb{R}^{n}$.
Theorem 3.1.
Let $n\geq 2$ and $0\leq\alpha<\alpha_{n}.$ Denote
$AT\left(\alpha,\beta\right)=\sup_{\left\|\nabla u\right\|_{n}\leq 1}\frac{1}{\left\|u\right\|_{n}^{n-\beta}}\int_{\mathbb{R}^{n}}\phi_{n}\left(\alpha\left(1-\frac{\beta}{n}\right)\left|u\right|^{\frac{n}{n-1}}\right)\frac{dx}{\left|x\right|^{\beta}}.$
Then there exist positive constants $c=c\left(n,\beta\right)$ and $C=C\left(n,\beta\right)$ such that when $\alpha$ is close enough to $\alpha_{N}:$
(3.1) $\frac{c\left(n,\beta\right)}{\left(1-\left(\frac{\alpha}{\alpha_{n}}\right)^{n-1}\right)^{\left(n-\beta\right)/}}\leq AT\left(\alpha,\beta\right)\leq\frac{C\left(n,\beta\right)}{\left(1-\left(\frac{\alpha}{\alpha_{n}}\right)^{n-1}\right)^{\left(n-\beta\right)/n}}.$
Moreover, the constant $\alpha_{N}$ is sharp in the sense that $AT\left(\alpha_{N},\beta\right)=\infty.$
Theorem 3.2 (Uniform local estimate).
Given a complete and noncompact Riemannian manifold $M$. Assume there exist $\delta_{0}>0$ and $\Lambda\geq 1$ such that for every $p\in M$ there is a harmonic coordinate chart
$\varphi_{p}:B_{\delta_{0}}(p)\longrightarrow\mathbb{R}^{n}$
whose image contains a Euclidean ball $B_{r}(0)$ with $r\in[c_{1}\delta_{0},c_{2}\delta_{0}]$, and such that on $B_{\delta_{0}}(p)$
$\Lambda^{-1}\left\lvert\xi\right\rvert^{2}\leq g^{ij}(x)\xi_{i}\xi_{j}\leq\Lambda\left\lvert\xi\right\rvert^{2},\qquad\Lambda^{-1}\leq\sqrt{\det(g_{ij}(x))}\leq\Lambda.$
Then there exists $\delta\in(0,\delta_{0}]$ and $C<\infty$, depending only on $(n,\tau,\Lambda)$, such that for every $u\in W^{1,n}(M)$ with $\left\lVert u\right\rVert_{1,\tau}\leq 1$,
$\sup_{p\in M}\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C.$
Proof.
Fix $\delta\in(0,\delta_{0}]$ (chosen at the end depending only on $(n,\tau,\Lambda)$) and fix $p\in M$. Let $\varphi:=\varphi_{p}$ be the harmonic chart on $B_{\delta}(p)$. Write $B_{r}:=B_{r}(0)\subset\varphi(B_{\delta}(p))$ with $r\sim\delta$ (constants depending only on $\Lambda$), and define
$\widetilde{u}(x):=u(\varphi^{-1}(x)),\qquad x\in B_{r}.$
By the metric and Jacobian comparability in harmonic coordinates, Euclidean and Riemannian $L^{n}$-norms on $B_{\delta}(p)$ and $B_{r}$ are comparable up to constants depending only on $\Lambda$.
Step 1: Cutoff (localization). Choose $\eta\in C_{c}^{\infty}(B_{r})$ such that
$\eta\equiv 1\ \text{on }B_{r/2},\qquad 0\leq\eta\leq 1,\qquad\left\lvert\nabla\eta\right\rvert\leq\frac{C}{r}.$
For a scaling parameter $a>0$ (to be chosen), set
$v(x):=a\,\widetilde{u}(x)\quad(x\in B_{r}),\qquad w:=\eta v,$
and extend $w$ by 0 to $\mathbb{R}^{n}$. Then $w\in W^{1,n}(\mathbb{R}^{n})$ and $w=v$ on $B_{r/2}$, hence
(3.2) $\int_{B_{r/2}}\phi_{n}\!\left(\alpha_{n}\left\lvert v\right\rvert^{\frac{n}{n-1}}\right)\,dx=\int_{B_{r/2}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx.$
Step 2: Gradient control for the cutoff. Since $\nabla w=\eta\nabla v+v\nabla\eta$,
$\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq C\left(\left\lVert\nabla v\right\rVert_{L^{n}(B_{r})}+\frac{1}{r}\left\lVert v\right\rVert_{L^{n}(B_{r})}\right).$
Using $v=a\widetilde{u}$, $r\sim\delta$, and harmonic coordinate comparability, we obtain
(3.3) $\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq C_{1}a\left(\left\lVert\nabla u\right\rVert_{L^{n}(B_{\delta}(p))}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(B_{\delta}(p))}\right)\leq C_{1}a\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right).$
Step 3: Choose $a$ uniformly so that $\left\lVert\nabla w\right\rVert\leq\theta$. Assume $\left\lVert u\right\rVert_{1,\tau}\leq 1$, i.e.
$\tau\left\lVert u\right\rVert_{L^{n}(M)}+\left\lVert\nabla u\right\rVert_{L^{n}(M)}\leq 1.$
Fix a margin $\theta:=\frac{1}{2}$ and choose
(3.4) $a:=\frac{1}{2C_{1}\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right)}.$
Then (3.3) gives
(3.5) $\left\lVert\nabla w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq\theta=\frac{1}{2}.$
(Here $a$ depends on $u$, but the final bound will be uniform because $a^{n}\left\lVert u\right\rVert_{L^{n}}^{n}$ is uniformly controlled by $\left\lVert u\right\rVert_{1,\tau}\leq 1$.)
Step 4: Apply the asymptotic estimates for the subcritical Trudinger-Moser inequality at a fixed subcritical parameter. Define
[…]
Then
$\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}=\alpha_{n}\theta^{\frac{n}{n-1}}\left\lvert z\right\rvert^{\frac{n}{n-1}}=:\alpha\,\left\lvert z\right\rvert^{\frac{n}{n-1}},\qquad\alpha:=\alpha_{n}\theta^{\frac{n}{n-1}}<\alpha_{n},$
where $\alpha$ is a fixed subcritical constant depending only on $n$. By the definition of $\mathrm{AT}(\alpha,0)$,
(3.6) $\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx=\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha\left\lvert z\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq\mathrm{AT}(\alpha,0)\,\left\lVert z\right\rVert_{L^{n}(\mathbb{R}^{n})}^{n}=\mathrm{AT}(\alpha,0)\,\theta^{-n}\,\left\lVert w\right\rVert_{L^{n}(\mathbb{R}^{n})}^{n}.$
Since $\alpha$ is fixed and $\alpha<\alpha_{n}$, $\mathrm{AT}(\alpha,0)$ is a finite constant depending only on $n$.
Step 5: Uniform control of $\left\lVert w\right\rVert_{L^{n}}$ and conclusion.
Because $0\leq\eta\leq 1$ and $w=\eta v$ on $B_{r}$,
$\left\lVert w\right\rVert_{L^{n}(\mathbb{R}^{n})}\leq\left\lVert v\right\rVert_{L^{n}(B_{r})}=a\,\left\lVert\widetilde{u}\right\rVert_{L^{n}(B_{r})}\leq C_{2}\,a\,\left\lVert u\right\rVert_{L^{n}(B_{\delta}(p))}\leq C_{2}\,a\,\left\lVert u\right\rVert_{L^{n}(M)}.$
Substituting into (3.6) yields
$\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq C_{3}(n)\,a^{n}\,\left\lVert u\right\rVert_{L^{n}(M)}^{n}.$
Now use the definition (3.4).
Since
$\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\geq\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)},$
we obtain
$a=\frac{1}{2C_{1}\left(\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{1}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}\right)}\leq\frac{\delta}{2C_{1}}\,\frac{1}{\left\lVert u\right\rVert_{L^{n}(M)}}.$
Therefore
$a^{n}\,\left\lVert u\right\rVert_{L^{n}(M)}^{n}\leq\left(\frac{\delta}{2C_{1}}\right)^{n},$
and we conclude
$\int_{\mathbb{R}^{n}}\phi_{n}\!\left(\alpha_{n}\left\lvert w\right\rvert^{\frac{n}{n-1}}\right)\,dx\leq C_{4}(n,\Lambda)\,\delta^{n}.$
Combining with (3.2) gives the Euclidean local bound on $B_{r/2}$. Finally, pulling back by $\varphi$ and using Jacobian comparability yields
$\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C(n,\tau,\Lambda),$
with $C$ independent of $u$ and $p$ (after fixing $\delta$ depending only on $(n,\tau,\Lambda)$). This proves the theorem. ∎
Next we need to putting together the local estimates over balls to get the estimates over $M$.
Lemma 3.1 (Cutoff bounds).
Let $\{x_{j}\}$ be as in Lemma 2.2. For each $j$ choose $\psi_{j}\in C_{c}^{\infty}(B_{\delta}(x_{j}))$ such that $0\leq\psi_{j}\leq 1$, $\psi_{j}\equiv 1$ on $B_{\delta/2}(x_{j})$, and $\left\lvert\nabla\psi_{j}\right\rvert\leq 4/\delta$. Then $\left\lvert\nabla(\psi_{j}^{2})\right\rvert\leq 8\psi_{j}/\delta$ and for every $u\in W^{1,n}(M)$ and every $\widetilde{\tau}\geq 0$,
(3.7) $\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\Bigl(\widetilde{\tau}+\frac{8}{\delta}\Bigr)\left\lVert u\right\rVert_{L^{n}(M)}.$
Moreover,
(3.8) $\sum_{j}\int_{M}\psi_{j}^{2n}\left\lvert\nabla u\right\rvert^{n}\,dV_{g}\leq N\int_{M}\left\lvert\nabla u\right\rvert^{n}\,dV_{g},\qquad\sum_{j}\int_{M}\left\lvert\nabla(\psi_{j}^{2})\right\rvert^{n}\left\lvert u\right\rvert^{n}\,dV_{g}\leq\frac{8^{n}}{\delta^{n}}\,N\int_{M}\left\lvert u\right\rvert^{n}\,dV_{g}.$
Proof.
For (3.7), compute
$\nabla(\psi_{j}^{2}u)=\psi_{j}^{2}\nabla u+u\,\nabla(\psi_{j}^{2}),$
hence by the triangle inequality in $L^{n}$,
$\left\lVert\nabla(\psi_{j}^{2}u)\right\rVert_{L^{n}(M)}\leq\left\lVert\psi_{j}^{2}\nabla u\right\rVert_{L^{n}(M)}+\left\lVert u\,\nabla(\psi_{j}^{2})\right\rVert_{L^{n}(M)}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{8}{\delta}\left\lVert\psi_{j}u\right\rVert_{L^{n}(M)}\leq\left\lVert\nabla u\right\rVert_{L^{n}(M)}+\frac{8}{\delta}\left\lVert u\right\rVert_{L^{n}(M)}.$
Also $\left\lVert\psi_{j}^{2}u\right\rVert_{L^{n}}\leq\left\lVert u\right\rVert_{L^{n}}$, so adding $\widetilde{\tau}\left\lVert\psi_{j}^{2}u\right\rVert_{L^{n}}$ gives (3.7). The implication
$\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq 1$
follows immediately from $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ and $\left\lVert u\right\rVert_{1,\tau}\leq 1$.
Theorem 3.3 (Global Moser–Trudinger from the local estimate).
2.2 Assume the hypotheses of Lemma  2.2. Fix $\delta\in(0,r_{H}]$ and assume the following local estimate holds:
(Local MT at scale $\delta$). There exist $\widetilde{\tau}\geq 0$ and $C_{\mathrm{loc}}>0$ (depending only on $n$ and the geometric data) such that for every $p\in M$ and every $f\in W^{1,n}(M)$ supported in $B_{\delta}(p)$ with $\left\lVert f\right\rVert_{1,\widetilde{\tau}}\leq 1$,
(3.9) $\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert f\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{loc}}.$
Then for every $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ there exists $C_{\mathrm{glob}}<\infty$ depending only on $(n,\tau)$ and the geometric data such that for all $u\in W^{1,n}(M)$ with $\left\lVert u\right\rVert_{1,\tau}\leq 1$,
(3.10) $\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{glob}}.$
Proof.
Let $\{x_{j}\}$ and $\{\psi_{j}\}$ be as in Lemmas  2.2 and 3.1. Since $M=\bigcup_{j}B_{\delta/2}(x_{j})$ and $\psi_{j}\equiv 1$ on $B_{\delta/2}(x_{j})$, we have pointwise on $B_{\delta/2}(x_{j})$:
$\left\lvert u\right\rvert=\left\lvert\psi_{j}^{2}u\right\rvert$.
Using that $\phi_{n}$ is increasing on $0,\infty)$, we obtain
(3.11) $\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}=\sum_{j}\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert\psi_{j}^{2}u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}.$
Apply the local estimate with correct scaling. Assume $\tau\geq\widetilde{\tau}+\frac{8}{\delta}$ and $\left\lVert u\right\rVert_{1,\tau}\leq 1$. Then Lemma 3.1 gives $\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}\leq 1$ for all $j$, and $\psi_{j}^{2}u$ is supported in $B_{\delta}(x_{j})$. Thus we may apply (3.9) with $p=x_{j}$ and $f=\psi_{j}^{2}u$ to obtain
$\int_{B_{\delta/2}(x_{j})}\phi_{n}\!\left(\alpha_{n}\left\lvert\psi_{j}^{2}u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C_{\mathrm{loc}}.$
To globalize, we use the standard homogeneous consequence of the local estimate:
Homogeneous local bound. There exists $C^{\prime}_{\mathrm{loc}}>0$ such that for all $p\in M$ and all $f$ supported in $B_{\delta}(p)$ and satisfying $\left\lVert f\right\rVert_{1,\widetilde{\tau}}\leq 1$,
(3.12) $\int_{B_{\delta/2}(p)}\phi_{n}\!\left(\alpha_{n}\left\lvert f\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\;\ \left\lVert f\right\rVert_{1,\widetilde{\tau}}^{\,n}.$
Applying (3.12) to $f=\psi_{j}^{2}u$ and summing in (3.11) gives
$\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\sum_{j}\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}.$
Estimate the sum of localized norms using bounded overlap. Using
$(A+B)^{n}\leq 2^{n-1}(A^{n}+B^{n}),$
the product rule, and $\psi_{j}\leq 1$,
$\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}\leq C(n)\int_{M}\psi_{j}^{2n}\left\lvert\nabla u\right\rvert^{n}+\left\lvert\nabla(\psi_{j}^{2})\right\rvert^{n}\left\lvert u\right\rvert^{n}+\widetilde{\tau}^{n}\psi_{j}^{2n}\left\lvert u\right\rvert^{n}\,dV_{g}.$
Summing over $j$ and using Lemma 3.1 and $\sum_{j}\psi_{j}^{2n}\leq N$) yields
$\sum_{j}\left\lVert\psi_{j}^{2}u\right\rVert_{1,\widetilde{\tau}}^{\,n}\leq C(n)N\int_{M}\left\lvert\nabla u\right\rvert^{n}\,dV_{g}+C(n)N\left(\frac{8^{n}}{\delta^{n}}+\widetilde{\tau}^{n}\right)\int_{M}\left\lvert u\right\rvert^{n}\,dV_{g}.$
If $\left\lVert u\right\rVert_{1,\tau}\leq 1$ then $\int_{M}\left\lvert\nabla u\right\rvert^{n}\leq 1$ and $\int_{M}\left\lvert u\right\rvert^{n}\leq\tau^{-n}$. Therefore
$\int_{M}\phi_{n}\!\left(\alpha_{n}\left\lvert u\right\rvert^{\frac{n}{n-1}}\right)\,dV_{g}\leq C^{\prime}_{\mathrm{loc}}\,C(n)N\left(1+\left(\frac{8^{n}}{\delta^{n}}+\widetilde{\tau}^{n}\right)\tau^{-n}\right)=:C_{\mathrm{glob}},$
which is finite and depends only on $(n,\tau)$ and the geometric data. This proves (3.10). ∎
4. Green’s function estimate: Proof of Theorem 1.2
@@
where $H(x,y)=\frac{1}{(n-2)\omega_{n-1}}d(x,y)^{2-n}f(d(x,y))$ is the fundamental solution of Laplacian multiplied by a smooth cutoff function $f$, which is identically equal to 1 when $d(x,y)<\frac{\delta}{2}$ and vanishes when $d(x,y)\geq\delta$. Moreover, $f$ can be chosen such that $|f^{\prime}|\leq\frac{C}{\delta},|f^{\prime\prime}|\leq\frac{C}{\delta^{2}}$ for some uniform constant $C>0$. $\Gamma(x,y)=\Gamma_{1}(x,y)=-\Delta_{y}H(x,y)$ and we inductively define
@@
where $C=C(n,\lambda,Q,\delta)$. Indeed, from the definition of $\Gamma_{l}(x,y)$, their singularity decrease as $l$ grows hence we first look at
@@
For the asymptotic behavior near singularity, we need to prove the following lemma:
@@
(4.4) $=$ $\left(\int_{B_{\rho}(x)\cap B_{\delta}(x_{j})}+\int_{(B_{3\rho}(y)\setminus B_{\rho}(x))\cap B_{\delta}(x_{j})}+\int_{B_{\delta}(x_{j})\setminus B_{3\rho}(y)}\right)d(x,z)^{\alpha-n}d(z,y)^{\beta-n}dV_{g}(z).$
The first term of the right hand side can be estimated as
@@
$\leq$ $C(\lambda,\delta,Q)\omega_{n-1}\rho^{\beta-n}\int_{0}^{\rho}r^{\alpha-n}r^{n-1}dr,$
$\leq$ $C(\lambda,\delta,Q)\omega_{n-1}\rho^{\alpha+\beta-n},$
where from the second line to the third line we use the property that in harmonic coordinate, $Q^{-1}\delta_{ij}\leq g_{ij}\leq Q\delta_{ij}$ as bilinear form. Hence integral can be transferred to normal coordinate and further to polar coordinate. Meanwhile it is worth to emphasize that we omit the transition function between harmonic coordinate charts centered at $x_{j}$ and centered at $x$, since it is a $C^{1,\alpha}$ function and can just be treated as a uniform constant. We will also apply such treatment in future estimates. Similarly we can estimate the second term of the right hand side of (4.3). For the third term, we have
@@
(4.7) $\leq$ $C(n,\lambda,\delta,q)\int_{3\rho}^{\delta}r^{\alpha+\beta-2n}r^{n-1}dr,$